A 30 g marble rolls with a velocity of 2 m/s toward a 70 g marble at rest. The lighter marble (m1) stops moving as its hits the heavier marble (m2). This collision causes the heavier marble to move. What is the velocity of the heavier marble?


1
Expert's answer
2021-11-07T19:27:42-0500

We can find the velocity of the heavier marble after the collision from the law of conservation of momentum:


"m_1v_{1i}+m_2v_{2i}=m_1v_{1f}+m_2v_{2f},""m_1v_{1i}=m_2v_{2f},""v_{2f}=\dfrac{m_1v_{1i}}{m_2}=\dfrac{0.03\ kg\times2\ \dfrac{m}{s}}{0.07\ kg}=0.86\ \dfrac{m}{s}."